OpenStudy (anonymous):

how do u find the slope of a line, give me a explanation a third grader can uunderstand

OpenStudy (anonymous):

slope (usually called "m") = change in y / change in x = ∆y/∆x. For instance, with the line y=x+1, the y value increases by 1 every time the x value increases by 1 - therefore, the slope is 1/1=1.

OpenStudy (anonymous):

OMGGGG THANK U!!!!

OpenStudy (anonymous):

now.... y intersept

OpenStudy (anonymous):

Y-intercept is basically any point that lies along the y-axis: which means, when x=0, what is the y value? Again, for the line y=x+1, the y-intercept is 1 because y=1 when x=0.